OpenStudy (anonymous):

Find the area of the triangle- B- 1 3/4 yards H- 4/5 yard A= 1/2( 1 3/4 x 4/5) ?

OpenStudy (anonymous):

A= 1/2bh

OpenStudy (anonymous):

\[1\frac{ 3 }{ 4 }=\frac{1* 4+3 }{ 4 }=\frac{ 7 }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }*\frac{ 7 }{ 4}*\frac{ 4 }{ 5 }=\frac{ 7 }{ 10 }=Area\]